A playground is on the flat roof of the city school 6m about the street. The ball takes 2.2 seconds to reach a point vertically above the wall.


a. Find initial velocity at which the ball was launch.


b. Find the vertical distance by which the ball clears the wall


1
Expert's answer
2021-05-19T12:43:04-0400

Assume that " The vertical wall of the building is h=7.00 m high, forming a 1 m high railing around the playground. A ball has fallen to the street below, and a passerby returns it by launching it at an angle of θ=53.00° above the horizontal at a point d=24.0 m from the base of the building wall".


a. "d=v_0\cos53°\cdot t\to v_0=d/(\cos53°\cdot t )=24/(\cos53°\cdot2.2)=18.1\ (m/s)" . Answer


b. "h=v_0\sin53°\cdot t -gt^2/2=18.1\cdot\sin53°\cdot2.2-9.8\cdot2.2^2/2=8.1\ (m)"


"8.1-7=1.1\ (m)" . Answer